OpenStudy (anonymous):

Solve the equation by completing the square,x^2+4x+3=0

OpenStudy (uri):

Quadratic equation?

OpenStudy (anonymous):

ya

OpenStudy (uri):

You know the formula?

Parth (parthkohli):

Why use the formula? I mean, you can factorize...

OpenStudy (uri):

Factorize then :D

OpenStudy (aravindg):

@uri "Solve the equation by completing the square"

OpenStudy (anonymous):

im so confused i asked for help but idk how to do it

OpenStudy (uri):

I know this wait :P

OpenStudy (anonymous):

factorize doesnt work

OpenStudy (anonymous):

okay thanks

Parth (parthkohli):

It does.

OpenStudy (uri):

x^2+4x+3=0 x^2+1x+3x+3 =0 x(x+1)+3(x+1) =0 (x+1) (x+3) =0 Divide them into 2 parts x+1=0====>x=-1 x+3=0====>x=-3

OpenStudy (anonymous):

do i say x= or by self

OpenStudy (uri):

Wut.

OpenStudy (anonymous):

x=-1,3 pr 1,-3

OpenStudy (anonymous):

or*

OpenStudy (anonymous):